# LLM-3D Print: Large Language Models To Monitor and Control 3D Printing

### v1

@@ Abstract
Industry 4.0 has revolutionized manufacturing by driving digitalization and shifting the paradigm toward additive manufacturing (AM). Fused Deposition Modeling (FDM), a key AM technology, enables the creation of highly customized, cost-effective products with minimal material waste through layer-by-layer extrusion, posing a significant challenge to traditional subtractive methods. However, the susceptibility of material extrusion techniques to errors often requires expert intervention to detect and mitigate defects that can severely compromise product quality. While automated error detection and machine learning models exist, their generalizability across diverse 3D printer setups, firmware, and sensors is limited, and deep learning methods require extensive labeled datasets, hindering scalability and adaptability. To address these challenges, we present a process monitoring and control framework that leverages pre-trained Large Language Models (LLMs) alongside 3D printers to detect and address printing defects. The LLM evaluates print quality by analyzing images captured after each layer or print segment, identifying failure modes and querying the printer for relevant parameters. It then generates and executes a corrective action plan. We validated the effectiveness of the proposed framework in identifying defects by comparing it against a control group of engineers with diverse AM expertise. Our evaluation demonstrated that LLM-based agents not only accurately identify common 3D printing errors, such as inconsistent extrusion, stringing, warping, and layer adhesion, but also effectively determine the parameters causing these failures and autonomously correct them without any need for human intervention.
@@ 1 Introduction
Some notable approaches to address these problems involve simulating the 3D printing process and optimizing process parameters specific to each part 34, 30. While this method can effectively tailor parameters for individual parts, it is computationally expensive and impractical for manufacturing due to the diverse range of parts and the significant time required for simulations. Another approach to solving failures in FDM focuses on its open-loop control system, which lacks real-time feedback and prevents immediate detection and correction of errors. By converting to a closed-loop system, failures can be detected, and print parameters optimized in real-time, addressing issues related to machine variability, user settings, and part-agnostic print parameters 35. This limitation has spurred innovative research into monitoring and improving extrusion-based additive manufacturing methods 36, 37, 38, 39. Researchers have employed various sensors, including vibration sensors 40, acoustic sensors 41, accelerometers 42, piezo sensors 43, and laser scanners 44 to enhance the detection and correction of errors in real-time. While these sensors can detect real-time anomalies and reduce print failures, many errors remain undetectable, and their high cost and complexity limit their widespread use in 3D printers 45.
@@
In contrast, camera-based monitoring systems offer a low-cost and easy-to-integrate alternative for identifying surface defects and other visible issues in 3D printing 46, 47, 48, 49. Infrared or thermal cameras can detect anomalies not visible to standard visual cameras 50, 51, and multi-camera setups can reconstruct 3D models of printed parts to check dimensional and shape accuracy 52, 53. Due to their cost-effectiveness and simplicity, standard camera systems have seen more widespread adoption. A notable study developed an image-based closed-loop quality control system for fused filament fabrication, implemented by a customized online image acquisition system with a proposed image diagnosis-based feedback quality control method 54. While computer vision approaches are promising for targeting specific errors, they often require calibration for each part, printer, and material, making it difficult to create feature extraction algorithms that generalize across different setups. Consequently, these methods typically work with only a single combination of printer, part geometry, material, and printing conditions. Integrating machine learning techniques offers a more flexible and adaptable solution, enabling robust error detection and correction across diverse printing environments.
@@
Machine learning, particularly deep learning techniques, has achieved state-of-the-art performance in various science and engineering applications. These techniques have been successfully applied to surrogate modeling 55, 56, 57, 58, generating 3D printable structures 59, and predicting failures in 3D printing 60, 61, 62, among other areas 63, 64. These advancements demonstrate the significant potential of machine learning to enhance and innovate within the field of additive manufacturing.
@@
Recent research has leveraged machine learning for error detection in FDM, demonstrating promising results through camera-based monitoring and convolutional neural networks to predict and correct issues such as extrusion rate 65, warpage 66, surface defects 67, and layer defects 68. One notable approach aims to generalize the process across different parts, materials, and printing systems by using convolutional neural networks to detect errors and adjust parameters like flow rate, speed, z-offset, and extruder temperature 69.
@@
However, most methods are limited to addressing errors in a single modality by changing only a few print parameters and have primarily demonstrated correction of the flow rate parameter in a single geometry used for both training and testing. Additionally, some machine learning models require a reference object for comparison, which restricts their effectiveness for custom parts 67. Furthermore, these methods are not capable of real-time correction, meaning that if an error is detected, the part cannot be recovered. Although Brion et al. 69 addressed this by splitting the toolpath for a layer into smaller segments, optimizing one segment does not necessarily ensure that subsequent segments will also be optimized, especially if their shapes differ significantly.
@@
Leveraging transformer architecture 70, 71 and massive datasets 72, 73, 74, 75, 76, 77, Large Language Models (LLMs) have made significant strides in various natural language processing (NLP) tasks, including text generation and following task-specific instructions 78, 79. Additionally, LLMs have demonstrated emergent reasoning capabilities by interpolating and utilizing their extensive training data, allowing them to make inferences, draw conclusions, and solve problems beyond their explicit programming 80, 81. Despite the substantial computational resources required for training and fine-tuning LLMs for specific applications, these models have shown an exceptional ability to generalize to new tasks and domains. This remarkable capability is primarily attributed to the in-context learning (ICL) paradigm 72. By using minimal natural language prompts and avoiding extensive fine-tuning, LLMs have emerged as effective ”few-shot learners” 82, 83, demonstrating proficiency with limited training examples.
@@
The adaptability of Large Language Models (LLMs) to new domains and their ability to learn in context has been transformative across numerous scientific fields. In chemistry, for example, LLMs have autonomously designed, planned, and executed complex experiments 84, 85. In mathematics and computer science, these models have discovered novel solutions to longstanding problems like the cap set problem and optimized algorithms for challenges such as the bin-packing problem 86. Furthermore, LLMs have significantly advanced research in biomedical fields 87, 88, 89, materials science 90, 91, and environmental science 92. Their impact extends to other scientific domains as well, enhancing understanding and expanding capabilities 93, 94, 95, 96, 97. Additionally, LLMs have proven to be effective optimizers for foundational problems such as linear regression and the traveling salesman problem, often matching or exceeding the performance of specialized heuristics through straightforward prompting 98. This versatility and efficacy underscore the potential of LLMs to drive innovation and efficiency in a wide array of scientific and engineering applications.
@@
In the field of mechanical engineering, fine-tuned Large Language Models (LLMs) have demonstrated exceptional capabilities in advanced tasks such as knowledge retrieval, hypothesis generation, and agent-based modeling. They have played a crucial role in integrating diverse domains through the use of knowledge graphs 99, 100. Additionally, LLMs excel in various design-related tasks, including sketch similarity analysis, material selection, engineering drawing analysis, CAD generation, and structural optimization 101, 102. Their ability to seamlessly connect and enhance these tasks highlights their transformative potential in mechanical engineering.
@@ 2.1 Error Detection
Large Vision-Language Models (LVLMs) such as MiniGPT-4 103 and LLaVA 104 have demonstrated significant capabilities in understanding and interpreting images, achieving remarkable performance across various visual tasks 105, 106, 107, 108. Owing to extensive pre-training on large datasets of image-text pairs, these models possess exceptional generalization abilities, enabling impressive zero-shot performance. They can effectively recognize and process both familiar and unfamiliar objects without the need for additional training data. This capability to generalize from diverse datasets makes them highly versatile and powerful tools for a wide range of visual and multimodal applications.
In this work, we leverage GPT-4o’s 77 image analysis capabilities to detect anomalies in ongoing 3D prints, enhancing the accuracy and reliability of the printing process by identifying and addressing issues in near real-time. After each layer is printed, the printer pauses, and two frame-mounted cameras capture images—one from the top (extruder view) and one from the front—after the extruder returns to its home position. These images along with a description of the part are fed into the LLM for analysis. To prevent redundant error identification, images from the last printed layer are also provided to the LLM, ensuring it does not detect previously addressed issues.
@@ 2.3 Planning Agent
While LLMs have demonstrated emergent reasoning capabilities, their performance can be significantly enhanced when guided by a structured reasoning framework 109, 110. Additionally, because the output of an LLM is heavily influenced by the quality and specificity of the prompts, it becomes essential to adapt these prompts based on the specific problems being addressed. Providing a well-defined reasoning structure and carefully tailored prompts allows the LLM to deliver more accurate and effective solutions, ensuring better outcomes for complex tasks111.
@@ 2.4 Executor Agent
The Agent Executor (Fig. 3c) , a crucial component of the framework, utilizes the ReAct (Reasoning and Acting) method 112 to execute the generated plans.
@@ 4 Result
For the multi-layer prints, we conducted two primary experiments: one involving the printing of a spanner and the other featuring raised text, designed to simulate the complexities of multi-part prints. In these experiments, the optimization was done after the completion of each layer, allowing us to continuously monitor and optimize the printing process throughout the entire build.
@@
Additionally, to evaluate the framework’s adaptability to different materials and its ability to optimize parameters, we conducted tests using two different materials: PLA and TPU. Both prints were performed at a temperature of 190 °C, with the only variation being the print speed, which was set to 120 mm/s, and a layer height of 0.35mm. These tests aimed to verify the framework’s capability to work effectively across different materials and optimize parameters accordingly.
@@ 4.1 Multi Layer Prints
Additionally, for the print of the wrench model, a manual perturbation was introduced at Layer 10 to disrupt the correct Z-axis movement of the extruder, causing it to shift randomly. This was done to evaluate whether the LLM could accurately detect and respond to sudden, unexpected changes in the printing process. This disruption was not implemented programmatically through the printer’s API, as the LLM might have simply traced the last issued commands and recognized any intentional modifications to the Z-offset. By introducing a manual obstruction, the test aimed to challenge the LLM’s ability to identify and correct unplanned anomalies.
@@ 4.1.1 Parameter Optimization
4.2 Error Detection
@@
Figure 6: Failures detected (a) Observations and failures detected by the LLM in single-layer print during print optimization and for (b) Multi-Layer print of wrench. (b)(1) Comparison of collected human response and LLM response to multilayer print for each layer and (b)(2) the agreement score.
Error detection is a vital element of the proposed framework. While LLMs are highly effective at analyzing textual data and sensor logs due to their advanced natural language processing capabilities, identifying subtle errors in a 3D print is challenging, especially when the model has not been specifically trained or fine-tuned on 3D printing data. To address this, we analyze the observations and errors detected by the LLM during the printing of a single layer.
The LLM accurately identified several key issues (Fig 6 a )at each segment of the print. It detected under-extrusion, nozzle clogs, inconsistent printing speeds, and bed leveling problems. These identifications were made by analyzing the visual characteristics of each segment’s print, such as gaps, stringing, and uneven filament deposition. From these observations, the LLM drew inferences about potential causes and provided recommendations for addressing the detected issues.
To demonstrate the effectiveness and error detection capability of the LLM (GPT-4o) in multi-layer prints, we conducted a comprehensive survey involving a diverse group of engineers, including graduate students with varying levels of experience in Additive Manufacturing (AM). The participants in this survey ranged from individuals with basic exposure to AM to those with advanced expertise in the field. Each participant was tasked with evaluating multi-layer 3D prints, specifically focusing on the prints that the LLM had analyzed and optimized during the printing process.

### v2

@@ Abstract
Industry 4.0 has revolutionized manufacturing by driving digitalization and shifting the paradigm toward additive manufacturing (AM). Fused Deposition Modeling (FDM), a key AM technology, enables the creation of highly customized, cost-effective products with minimal material waste through layer-by-layer extrusion, posing a significant challenge to traditional subtractive methods. However, the susceptibility of material extrusion techniques to errors often requires expert intervention to detect and mitigate defects that can severely compromise product quality. While automated error detection and machine learning models exist, their generalizability across diverse 3D printer setups, firmware, and sensors is limited, and deep learning methods require extensive labeled datasets, hindering scalability and adaptability. To address these challenges, we present a process monitoring and control framework that leverages Large Language Models (LLMs) alongside 3D printers to detect and address printing defects—entirely without domain-specific fine-tuning or training. Instead, the LLM operates using in-context learning and general reasoning to interpret print quality from images captured after each layer or print segment, identify failure modes, and query the printer for relevant parameters. It then autonomously generates and executes a corrective action plan. We validated the effectiveness of the proposed framework in identifying defects by comparing it against a control group of engineers with diverse AM expertise. Our evaluation demonstrated that LLM-based agents not only accurately identify common 3D printing errors—such as inconsistent extrusion, stringing, warping, and layer adhesion—but also effectively determine the underlying causes and autonomously correct them without any need for human intervention.
@@ 1 Introduction
Some approaches address print failures by simulating the 3D printing process and optimizing part-specific parameters 34, 30. While effective, these simulations are computationally expensive and impractical for production, especially when compared to the lower inference cost/time of LLMs, which operate on a per-layer basis in seconds rather than hours. Another strategy targets the limitations of open-loop FDM systems by introducing closed-loop control, where “real-time” refers to parameter adjustments made at the end of each print layer or even a more finer sampling rate. Prior work has incorporated sensors such as accelerometers 35, acoustic sensors 36, and laser scanners 37 to detect anomalies. However, many defects, such as stringing or subtle layer separation, lack strong physical signatures and remain difficult to detect with such sensors. Additionally, the cost and complexity of these sensor systems hinder their widespread use 38, motivating more generalizable, multimodal approaches like the one proposed here.
@@
In contrast, camera-based monitoring systems offer a low-cost and easy-to-integrate alternative for identifying surface defects and other visible issues in 3D printing 39, 40, 41, 42. Infrared or thermal cameras can detect anomalies not visible to standard visual cameras 43, 44, and multi-camera setups can reconstruct 3D models of printed parts to check dimensional and shape accuracy 45, 46. Due to their cost-effectiveness and simplicity, standard camera systems have seen more widespread adoption. A notable study developed an image-based closed-loop quality control system for fused filament fabrication, implemented by a customized online image acquisition system with a proposed image diagnosis-based feedback quality control method 47. While computer vision approaches are promising for targeting specific errors, they often require calibration for each part, printer, and material, making it difficult to create feature extraction algorithms that generalize across different setups. Consequently, these methods typically work with only a single combination of printer, part geometry, material, and printing conditions. Integrating machine learning techniques offers a more flexible and adaptable solution, enabling robust error detection and correction across diverse printing environments.
@@
Machine learning, particularly deep learning techniques, has achieved state-of-the-art performance in various science and engineering applications. These techniques have been successfully applied to surrogate modeling 48, 49, 50, 51, generating 3D printable structures 52, and predicting failures in 3D printing 53, 54, 55, among other areas 56, 57. These advancements demonstrate the significant potential of machine learning to enhance and innovate within the field of additive manufacturing.
@@
Recent research has leveraged machine learning for error detection in FDM, demonstrating promising results through camera-based monitoring and convolutional neural networks to predict and correct issues such as extrusion rate 58, warpage 59, surface defects 60, and layer defects 61. One notable approach aims to generalize the process across different parts, materials, and printing systems by using convolutional neural networks to detect errors and adjust parameters like flow rate, speed, z-offset, and extruder temperature 62.
@@
However, most methods are limited to addressing errors in a single modality by changing only a few print parameters and have primarily demonstrated correction of the flow rate parameter in a single geometry used for both training and testing. Additionally, some machine learning models require a reference object for comparison, which restricts their effectiveness for custom parts 60. Furthermore, these methods are not capable of real-time correction, meaning that if an error is detected, the part cannot be recovered. Although Brion et al. 62 addressed this by splitting the toolpath for a layer into smaller segments, optimizing one segment does not necessarily ensure that subsequent segments will also be optimized, especially if their shapes differ significantly.
@@
Leveraging transformer architecture 63, 64 and massive datasets 65, 66, 67, 68, 69, 70, Large Language Models (LLMs) have made significant strides in various natural language processing (NLP) tasks, including text generation and following task-specific instructions 71, 72. Additionally, LLMs have demonstrated emergent reasoning capabilities by interpolating and utilizing their extensive training data, allowing them to make inferences, draw conclusions, and solve problems beyond their explicit programming 73, 74. Despite the substantial computational resources required for training and fine-tuning LLMs for specific applications, these models have shown an exceptional ability to generalize to new tasks and domains. This remarkable capability is primarily attributed to the in-context learning (ICL) paradigm 65. By using minimal natural language prompts and avoiding extensive fine-tuning, LLMs have emerged as effective ”few-shot learners” 75, 76, demonstrating proficiency with limited training examples.
@@
The adaptability of Large Language Models (LLMs) to new domains and their ability to learn in context has been transformative across numerous scientific fields. In chemistry, for example, LLMs have autonomously designed, planned, and executed complex experiments 77, 78. In mathematics and computer science, these models have discovered novel solutions to longstanding problems like the cap set problem and optimized algorithms for challenges such as the bin-packing problem 79. Furthermore, LLMs have significantly advanced research in biomedical fields 80, 81, 82, materials science 83, 84, and environmental science 85. Their impact extends to other scientific domains as well, enhancing understanding and expanding capabilities 86, 87, 88, 89, 90. Additionally, LLMs have proven to be effective optimizers for foundational problems such as linear regression and the traveling salesman problem, often matching or exceeding the performance of specialized heuristics through straightforward prompting 91. This versatility and efficacy underscore the potential of LLMs to drive innovation and efficiency in a wide array of scientific and engineering applications.
@@
In the field of mechanical engineering, fine-tuned Large Language Models (LLMs) have demonstrated exceptional capabilities in advanced tasks such as knowledge retrieval, hypothesis generation, and agent-based modeling. They have played a crucial role in integrating diverse domains through the use of knowledge graphs 92, 93. Additionally, LLMs excel in various design-related tasks, including sketch similarity analysis, material selection, engineering drawing analysis, CAD generation, and structural optimization 94, 95. Their ability to seamlessly connect and enhance these tasks highlights their transformative potential in mechanical engineering.
Figure 2: Schematic of the Proposed Framework. The process begins with a G-code file (a) being uploaded to the 3D printer, (b) which is equipped with two frame-mounted cameras. After each layer is printed, the extruder moves to the home position, and two images of the current print state are captured (c). These images are analyzed by the LLM, which evaluates the print, makes observations, and identifies any failures. If failures are detected, the LLM supervisor (d) invokes the information planner (e). The executor (f) then carries out the information gathering plan, after which the solution planner (g) is activated by the supervisor. The solution plan is executed by another executor (h), and finally, the supervisor invokes the handoff module (i) to resume the print.
@@ 2.1 Error Detection
Large Vision-Language Models (LVLMs) such as MiniGPT-4 96 and LLaVA 97 have demonstrated significant capabilities in understanding and interpreting images, achieving remarkable performance across various visual tasks 98, 99, 100, 101. Owing to extensive pre-training on large datasets of image-text pairs, these models possess exceptional generalization abilities, enabling impressive zero-shot performance. They can effectively recognize and process both familiar and unfamiliar objects without the need for additional training data. This capability to generalize from diverse datasets makes them highly versatile and powerful tools for a wide range of visual and multimodal applications.
In this work, we leverage GPT-4o’s 70 image analysis capabilities to detect anomalies in ongoing 3D prints, enhancing the accuracy and reliability of the printing process by identifying and addressing issues in near real-time. After each layer is printed, the printer pauses, and two frame-mounted cameras capture images—one from the top (extruder view) and one from the front—after the extruder returns to its home position. These images along with a description of the part are fed into the LLM for analysis. To prevent redundant error identification, images from the last printed layer are also provided to the LLM, ensuring it does not detect previously addressed issues.
@@ 2.3 Planning Agent
While LLMs have demonstrated emergent reasoning capabilities, their performance can be significantly enhanced when guided by a structured reasoning framework 102, 103. Additionally, because the output of an LLM is heavily influenced by the quality and specificity of the prompts, it becomes essential to adapt these prompts based on the specific problems being addressed. Providing a well-defined reasoning structure and carefully tailored prompts allows the LLM to deliver more accurate and effective solutions, ensuring better outcomes for complex tasks104.
@@ 2.4 Executor Agent
The Agent Executor (Fig. 3c) , a crucial component of the framework, utilizes the ReAct (Reasoning and Acting) method 105 to execute the generated plans.
@@ 3 Experimental Setup
In our evaluation, we used two different filament materials: PLA and TPU. To introduce environmental variability into the testing setup, the PLA filament had a thermochromic property that changes color from green to brown above 31°C, mimicking lighting inconsistencies or thermal gradients often encountered in uncontrolled environments. However, during the printing process, no noticeable color change was observed, suggesting that the thermal conditions remained stable throughout the print. As a result, the color-shifting property was not factored into the defect detection analysis. TPU was used in parallel to test the framework’s adaptability to more flexible and mechanically demanding materials.
To ensure robust and reliable interaction with the printer, the LLM was provided with a curated list of modifiable G-code objects derived from a summarized snapshot of the printer’s firmware. This approach was necessitated by early observations that supplying the full firmware documentation exceeded the model’s 128k token context limit, occasionally leading to unintended LLM-driven shutdowns—typically when the model misinterpreted persistent errors as indicators of critical system failure. To mitigate this, the curated summary excluded only sensitive or potentially destabilizing commands, such as those related to shutdown, restart, or permanent firmware modification. With these safeguards the LLM retained the ability to freely access and adjust a wide range of operational parameters essential to print quality, including fan speed, bed temperature, cruise speed, acceleration, etc. It could also dynamically query printer state via the API, maintaining its ability to reason over system conditions and implement corrective actions. This careful balance between safety and control enabled effective closed-loop operation without compromising hardware integrity. Additionally, we choose to ignore layer shift defect as such a defect is catastrophic to the print and it cannot be solved without discarding prints.
@@ 4 Result
For the multi-layer prints, we conducted two primary experiments: one involving the printing of a wrench and the other featuring raised text, designed to simulate the complexities of multi-part prints. In these experiments, the optimization was done after the completion of each layer, allowing us to continuously monitor and optimize the printing process throughout the entire build.
@@
Additionally, to evaluate the framework’s adaptability to different materials and its ability to optimize parameters, we conducted tests using two different materials: PLA and TPU. Both prints (single layer) were performed at a temperature of 190 °C, with the only variation being the print speed, which was set to 120 mm/s, and a layer height of 0.35mm. These tests aimed to verify the framework’s capability to work effectively across different materials and optimize parameters accordingly.
@@ 4.1 Multi Layer Prints
Additionally, for the print of the wrench model, a manual perturbation was introduced before Layer 9 to disrupt the correct Z-axis movement of the extruder, causing it to shift randomly. This was done to evaluate whether the LLM could accurately detect and respond to sudden, unexpected changes in the printing process. This disruption was not implemented programmatically through the printer’s API, as the LLM might have simply traced the last issued commands and recognized any intentional modifications to the Z-offset. By introducing a manual obstruction, the test aimed to challenge the LLM’s ability to identify and correct unplanned anomalies.
@@ 4.1.1 Parameter Optimization
5 Discussion
5.1 Analysis of LLM-generated Response
Figure 6: LLM-generated response at Layer 9. Expert-annotated image highlights observed defects: (1) extrusion inconsistency, (2) stringing, (3) layer separation, and (4) material blob. The corresponding LLM-generated response (formatted) identifies these issues and outlines corrective actions based on printer state and visual input.
As illustrated in Fig.6, the expert-annotated figure identifies four key failure modes: (1) extrusion inconsistency, (2) stringing, (3) layer separation, and (4) a material blob. The LLM successfully identified the first three issues and proposed actionable corrections. However, it did not flag the blob artifact, suggesting some limitations in recognizing less distinct defect types. The blob, formed by excess filament ending in fine strands, visually resembles stringing, which may have led to its misclassification. The model’s diagnostic reasoning was guided by domain-relevant questions generated through LLM-refined structured prompting (e.g., ”What could be causing stringing and poor adhesion?”) and did not rely on retraining or fine-tuning.
After identifying the defects, the LLM queried real-time printer parameters such as flow rate, retraction settings, fan speed, and Z-offset, then used this data to produce corrective G-code commands. It increased retraction length and speed to reduce stringing, adjusted the Z-offset for better adhesion, and raised the extrusion multiplier to address under-extrusion. This seamless flow from problem detection to intervention demonstrates the model’s ability to carry out structured, context-aware decisions.
The LLM also considered its previous corrective actions and adapted its strategy. It recognized that earlier adjustments had only partially improved the print and responded with stronger changes, increasing the flow rate from 105% to 110% and retraction speed from 50 to 55 mm/s. These changes were not arbitrary. The LLM linked stringing to retraction behavior, attributed poor adhesion to a manually introduced perturbation in the Z-offset at Layer 9, and associated under-extrusion with an insufficient flow rate. This demonstrates the ability to connect causes with the right process variables and propose meaningful adjustments.
By combining live printer data with structured reasoning, the LLM shows potential as an autonomous controller in manufacturing. It does not depend on fixed rules or pre-programmed workflows but instead adapts in real time using prompt-based logic. This makes it a promising tool for scalable and self-correcting quality control in additive manufacturing.
5.2 LLM Identified Defects
@@
Figure 7: Detected Failures. (a) Failures and observations identified by the LLM during single-layer print optimization. (b) Multi-layer print of the wrench model, showing a layer-by-layer comparison of detected failures across human annotators, domain experts, and the LLM.
Error detection is a vital component of the proposed closed-loop framework for additive manufacturing. While LLMs are highly effective at processing textual data and sensor logs, detecting visual or geometric print defects remains a significant challenge, particularly in the absence of fine-tuning on 3D printing-specific datasets. To assess this capability, we evaluated the LLM’s performance in identifying print defects during the fabrication of a single layer, comparing its outputs with expert annotations (SI: 8). In the absence of standardized metrics for print quality assessment, such comparisons provide a practical basis for evaluation. Prior work has employed annotated datasets and precision-recall analysis to assess model performance in similar contexts 38, 106, and we adopt a comparable approach.
In the single-layer prints, the LLM successfully identified several common failure modes, including stringing, under-extrusion, nozzle clogs, inconsistent print speeds, and bed leveling issues. These were inferred from visual indicators such as filament gaps, stray strands, and irregular deposition patterns (Fig. 7a). Based on these observations, the model proposed plausible root causes and recommended targeted corrective actions. The model’s outputs showed strong alignment with expert annotations, demonstrating its capability to perform real-time diagnostics on par with domain specialists, even without retraining.